An elementary school teacher is interested in knowing if there is a significant difference in the average reading speed of fifth-grade boys and girls. She randomly selects 36 fifth grades boys and 36 fifth grade girls for the study. Each student is given several pages of the same book to read and the time it takes them to complete the reading is recorded in minutes. Boys Girls Sample Size 36 36 Sample Mean reading speed (in minutes) 11 12 Population Variance 3 2 If you are interested in testing whether or not the average reading speed of boys is significantly quicker than that of girls, the test statistic is a. 1.5 b. 1.645 c. -1.75 d. -2.68
Answer #1

Pooled Variance
sp = sqrt(s1^2/n1 + s2^2/n2)
sp = sqrt(3.00017041/36 + 1.99996164/36)
sp = 0.3727

Test statistic,
z = (x1bar - x2bar)/sp
z = (11 - 12)/0.3727
z = -2.68